how would i make a linear equation out of this sentence: In a chemistry class, 12 L of a 12% alchol solution must be mixed with a 20% solution to get a 14% solution. How many liters of the 20% solution are needed?
Answer by stanbon(75887)

In a chemistry class, 12 L of a 12% alcohol solution must be mixed with a 20% solution to get a 14% solution. How many liters of the 20% solution are needed?
----
Equation:
alcohol + alcohol = alcohol
0.12(12) + 0.20x = 0.14(12+x)
-----------
Multiply thru by 100 to get:
12*12 + 20x = 14*12 + 14x
6x = 2*12
x = 4 liters (amt of 20% solution needed for the mixture)
----------------------
Cheers,
Stan H.
====================